# Weight Space Representation Learning on Diverse NeRF Architectures

### v2

@@ 4.3 NeRF Retrieval
The embeddings produced by the trained encoder can also be used to perform retrieval tasks via $k$-nearest neighbor search, as outlined in Fig. 1. In particular, we define instance-level retrieval as follows: given a NeRF embedding of a given object (a.k.a. the query) and a gallery of NeRF embeddings, the goal is to find the embedding in the gallery that represents the same object as the query but encodes a different NeRF architecture. We evaluate the performance on this task with the recall@$k$ metric [34], defined as the percentage of queries whose $k$ nearest neighbors contain the NeRF representing the same object as the query. Table 6 shows the retrieval results for various query/gallery test-set combinations. $\mathcal{L}_{\text{R}+\text{C}}$ outperforms $\mathcal{L}_{\text{R}}$ in all scenarios; in particular, when both query and gallery belong to ALL, $\mathcal{L}_{\text{R}}$ is unable to perform the task: the latent space organization shown in Fig. 3 (left) and discussed in Section 4.1 prevents $\mathcal{L}_{\text{R}}$ from being capable of recognizing the same object represented by NeRFs parameterized by different architectures. The slightly higher $\mathcal{L}_{\text{R}}$’s recall@$k$ in the other query/gallery combinations is due to the fact that the gallery is smaller and the chance of missing the target instance is thus lower. Conversely, the superior performance of $\mathcal{L}_{\text{R}+\text{C}}$ confirms that the contrastive objective favors a more architecture-agnostic latent space. Fig. 4 (top) displays qualitatively that the organization of the latent space learned by $\mathcal{L}_{\text{R}+\text{C}}$ captures the NeRF similarity in color and shape independently of the architecture.
Objaverse queries. To qualitatively assess the ability of our framework to generalize to unseen datasets, Fig. 4 (bottom) shows retrieval results when the gallery is the test set of MLP, TRI, or HASH, whereas the query belongs to a dataset of NeRFs trained on Objaverse [13], dubbed ObjaNeRF, recently released by Amaduzzi et al. [3]. ObjaNeRF’s NeRFs are vanilla MLPs with the same architecture as those in MLP (see Table 1). We point out that in this across-datasets experiment the query object is not present in the gallery. Remarkably, for queries that resemble objects featured in ShapenetRender [37], our method is able to retrieve NeRFs that belong to the same class as the query (rows 1–2 of Fig. 4, bottom), with some failure cases occurring with more unusual queries (row 3). An extended version of Fig. 4 is shown in Appendix E.
@@ 5 Conclusions
In this paper, we presented the first framework able to perform downstream tasks on NeRF weights with diverse architectures. Our method, based on a Graph Meta-Network encoder ingesting the NeRF parameter graph, is able to (i) handle input NeRFs parameterized by MLPs, tri-planes, and, for the first time, multi-resolution hash tables; (ii) process, at inference time, variations of architectures seen during training; (iii) provide state-of-the-art performance in the single architecture scenario. Furthermore, we investigated the interplay between rendering and contrastive training objectives and showed that they serve complementary purposes, by favoring either class-level separability or invariance to the NeRF architecture. The main limitation of our study lies in its experimental evaluation, which focuses on NeRFs that, although parameterized by 13 total architectures, are trained on a single dataset, i.e. ShapenetRender [37], with the notable exception of one generalization experiment on Objaverse [13]. Extending our training procedure and classification experiments on Objaverse itself or similar large-scale NeRF datasets would further validate the effectiveness of our approach, and we plan to follow this research path in future work. Eventually, our methodology could be scaled up to become the backbone to train a foundational model for NeRF weights processing.
@@ Appendix A Additional Parameter Graph Details
Figure 5: Parameter graph conversion. Top left: parameter graph representation of an MLP, proposed by Lim et al. [23]. Right: parameter graph representation of a tri-plane, proposed by Lim et al. [23]. Dotted edges should be connected to the $C$ channel nodes, but are not fully drawn for better visual clarity. Bottom left: our parameter graph representation of a multi-resolution hash table.
@@
The parameter graph conversion of an MLP, a tri-plane, and a multi-resolution hash table is depicted in Fig. 5, with additional details compared to Fig. 2 (left).
@@
MLP. The parameter graph of a linear layer (proposed by Lim et al. [23]) coincides with its computation graph, except for biases, which are modeled by including an additional node for each (non-input) layer, connected to every other neuron in the layer with an edge. As a result, the parameter graph of an MLP is identical to its computation graph, with the addition of the aforementioned bias nodes. Each node stores three integer values: layer index, neuron index within the layer, and node type (neuron or bias). Each edge stores the corresponding weight or bias and two integers, i.e. layer index and edge type (linear layer weight or linear layer bias).
@@
Tri-plane. The parameter graph of a tri-plane (proposed by Lim et al. [23]) contains one node for each spatial location and one node for each feature channel, where each spatial node is connected via an edge to every channel node; the channel nodes, together with one node for each input coordinate (more precisely, one node for each entry in their frequency encoding; see Appendix B) become the first layer of the MLP parameter graph that follows. Each node stores three integer values: layer index, neuron index within the layer, and node type (neuron for coordinate nodes or tri-planar for spatial and channel nodes). Each edge stores the corresponding learnable parameter (a.k.a. feature) and five additional values: layer index (integer), edge type (tri-planar, integer), and three $(i,j,k)$ values (evenly spaced floats $\in[-1,1]$) denoting the spatial position of the parameter within the tri-plane (with $i=0$ on the $yz$ plane, $j=0$ in the $xz$ plane, and $k=0$ on the $xy$ plane).
@@
Hash table. As mentioned in Section 3, we are the first to propose a parameter graph construction for multi-resolution hash tables. A single hash table is modeled with a node for each table entry and a node for each feature vector dimension, where each entry node is connected to each feature node via an edge. This subgraph construction is repeated for each hash table, i.e. for each resolution level. Finally, the concatenation of all feature nodes becomes the first layer of the MLP parameter graph that follows. Each node stores three integer values: layer index, neuron index within the layer, and node type (hash table). Each edge stores the corresponding learnable parameter (a.k.a. feature) and four additional integers: layer index, edge type (hash table), table index, and entry index within the table. Consider a multi-resolution hash table with $N$ levels, max table size $T$, $F$ features per entry, and max resolution $R$. Our parameter graph construction requires $N(T+F)$ nodes and $NTF$ edges. Conversely, if we drew an analogy with the tri-planar graph representation and explicitly modeled the underlying voxel grids with one node for each spatial location, it would result in a graph with $\mathcal{O}(N(R^{3}+F))$ nodes and $\mathcal{O}(NR^{3}F)$ edges. Since the purpose of hash grids is to reduce the memory footprint by mapping multiple voxel grid locations to the same table entry via a hashing function, $T$ should always be much smaller than $R^{3}$ (e.g. NeRFs in HASH have $T=2^{12}$ vs $R^{3}=2^{21}$), hence $N(T+F)\ll N(R^{3}+F)$ and $NTF\ll NR^{3}F$. Therefore, the motivation behind our choice of parameter graph representation for hash tables lies in its higher memory efficiency compared to its voxel-grid-based alternative.
Appendix B Additional Experimental Details
NeRF datasets. As outlined in Section 4, the datasets used throughout our experiments consist of NeRFs trained on ShapenetRender [37] (except for ObjaNeRF [3]). ShapenetRender is a dataset featuring RGB images and class labels of synthetic objects belonging to 13 classes of the ShapeNetCore dataset [9]: airplane, bench, cabinet, car, chair, display, lamp, speaker, rifle, sofa, table, phone, and watercraft. For each object, the dataset contains 36 $224\times 224$ images. We follow Zama Ramirez et al. [38]’s train-val-test split without augmentations, consisting of 31744 NeRFs used for training, 3961 for validation, and 3961 for testing. Table 7 summarizes the NeRF architectures featured in our experiments. In addition to the dataset descriptions provided in Section 4, we note here that:
  • In MLP and TRI, input $(x,y,z)$ coordinates go through a frequency encoding [26] before being concatenated to the tri-plane or hash table output that gets fed to the MLP, as this choice was made by the original creators of the two datasets, i.e. Zama Ramirez et al. [38] and Cardace et al. [7], respectively. Conversely, the MLP in HASH takes as input the output of the multi-resolution hash table by itself, without (encoded) input coordinates being concatenated to it; we choose to do so in order to be faithful to the original multi-resolution hash table architecture proposed by Mildenhall et al. [26]. The presence or absence of concatenated encoded coordinates in a dataset used for training is then reflected in all the unseen variations of that architecture family (e.g. MLP-2L uses frequency encoding, whereas HASH-2L does not).
  • MLP layers are intertwined with ReLU activation functions in MLP and HASH, whereas TRI uses sine activation functions [31]. The activation choice for MLP and TRI was made by their original creators [38, 7], whereas we choose ReLU activations for HASH in order to be faithful to the original hash table architecture [26]. The activation choice in a dataset used for training is then reflected in all the unseen variations of that architecture family (e.g. MLP-2L uses ReLU, whereas TRI-2L uses sine).
  • All NeRFs approximate a function from input coordinates to predicted color and density values, i.e. $(x,y,z)\mapsto(r,g,b,\sigma)$. This follows the simplified NeRF formulation proposed by Mildenhall et al. [26], the alternative being adding a viewing direction $(\theta,\phi)$ as input and letting the NeRF approximate a function $(x,y,z,\theta,\phi)\mapsto(r,g,b,\sigma)$. The formulation with viewing direction is usually modeled by two MLPs: a density MLP taking as input $(x,y,z)$ (or, more precisely, an encoding of $(x,y,z)$) and producing as output $\sigma$ together with an additional vector $\mathbf{h}$ of features, and a color MLP taking as input the concatenation of $\mathbf{h}$ and $(\theta,\phi)$ and predicting $(r,g,b)$ as output. Both Zama Ramirez et al. [38] and Cardace et al. [7] adopt the simplified $(x,y,z)\mapsto(r,g,b,\sigma)$ formulation modeled by a single MLP when creating their NeRF datasets (i.e. MLP and TRI, respectively). We choose to follow that same formulation for HASH and the datasets of unseen architectures, both to keep the datasets consistent and to reduce the complexity of the parameter graph going into the encoder.
@@ Appendix B Additional Experimental Details
Table 7: Dataset architectures. Extended version of Table 1. Architectural hyperparameters of NeRF datasets featured in our experiments, including those used at inference time only (unseen).
@@
MLP [38] TRI [7] HASH (ours) MLP-2L MLP-32H TRI-2L TRI-32H TRI-16W TRI-8C HASH-2L HASH-32H HASH-3N HASH-11T

### v3

@@ 4.3 NeRF Retrieval
The embeddings produced by the trained encoder can also be used to perform retrieval tasks via $k$-nearest neighbor search, as outlined in Fig.˜1. In particular, we define instance-level retrieval as follows: given a NeRF embedding of a given object (a.k.a. the query) and a gallery of NeRF embeddings, the goal is to find the embedding in the gallery that represents the same object as the query but encodes a different NeRF architecture. We evaluate the performance on this task with the recall@$k$ metric (Wang et al., 2017), defined as the percentage of queries whose $k$ nearest neighbors contain the NeRF representing the same object as the query. Tab.˜7 shows the retrieval results for various query/gallery test-set combinations; the recall@$k$ obtained by randomly shuffling the gallery is shown for reference (Random rows). Although both $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ significantly outperform the Random baseline, the superior performance of $\mathcal{L}_{\text{R}+\text{C}}$ confirms that the contrastive objective favors a more architecture-agnostic latent space. Indeed, Figs.˜6 and 7 display qualitatively that the latent space organization learned by $\mathcal{L}_{\text{R}+\text{C}}$ captures the NeRF similarity in color and shape independently of the underlying architecture.
Generalization on Objaverse. To assess the generalization capabilities of our framework, we leverage three additional datasets: $\texttt{MLP}^{\texttt{OB}}$, $\texttt{TRI}^{\texttt{OB}}$, and $\texttt{HASH}^{\texttt{OB}}$. With $\texttt{MLP}^{\texttt{OB}}$ we denote a dataset of MLP-based NeRFs trained on Objaverse (Deitke et al., 2023) recently released by Amaduzzi et al. (2025), whereas $\texttt{TRI}^{\texttt{OB}}$ and $\texttt{HASH}^{\texttt{OB}}$ are datasets, created by us, of tri-planar and hash-based NeRFs trained on the test split of Objaverse defined by Amaduzzi et al. Specifically, NeRFs in $\texttt{MLP}^{\texttt{OB}}$, $\texttt{TRI}^{\texttt{OB}}$, and $\texttt{HASH}^{\texttt{OB}}$ have the same architecture as those in MLP, TRI, and HASH, respectively (see Tab.˜1). Tab.˜7 mirrors the experiment in Tab.˜7, with the notable difference that the query and gallery contain NeRFs trained on a different dataset (i.e. Objaverse) than the one the model has seen during training (i.e. ShapenetRender (Xu et al., 2019)). Although this leads, unsurprisingly, to an overall drop in performance compared to Tab.˜7, it is worth noting, once again, that both $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ outperform the Random baseline, and the contrastive loss yields superior results. Fig.˜4 shows instead a qualitative retrieval experiment where the query is an MLP-based NeRF trained on Objaverse (i.e. $\texttt{MLP}^{\texttt{OB}}$) while the gallery contains NeRFs trained on ShapenetRender, hence the query object is not present in the gallery. Remarkably, for queries that resemble objects featured in ShapenetRender, our method retrieves NeRFs that belong to the same class as the query (rows 1–2 of Fig.˜4), with some failure cases occurring with more unusual queries (row 3). An extended version of these results is shown in Fig.˜8.
4.4 NeRF Captioning and Q&A
Amaduzzi et al. (2024; 2025) have previously shown that language tasks on NeRFs benefit from the holistic understanding of 3D objects captured by NeRF weights, leading their weight-processing framework (dubbed LLaNA) to outperform methods ingesting rendered views as input in NeRF captioning and NeRF Q&A tasks. To achieve these results, LLaNA uses the embeddings produced by the trained nf2vec encoder to condition an LLM belonging to the Llama 2 family (Touvron et al., 2023). To assess the effectiveness of our learned embeddings in downstream language tasks, we replaced the nf2vec encoder in LLaNA’s pipeline with our GMN encoder, trained the resulting framework on textual descriptions from the ShapeNeRF–Text dataset (Amaduzzi et al., 2024), and evaluated it on brief captioning, detailed captioning, and single-round Q&A. In the multi-architecture setting, our framework is trained on ALL (as usual), and the ALL embeddings produced by our encoder are used to train the LLM. Since the nf2vec encoder can process only MLP-based NeRFs, the original LLaNA framework cannot be applied in the multi-architecture setting. In the single-architecture setting, instead, our framework is trained on MLP, and the MLP embeddings produced by our encoder are used to train the LLM. In this setting, our results are compared with the original LLaNA framework, which is also trained on MLP. Multi-architecture results are shown in Tabs.˜13, 13 and 13, whereas single-architecture results are shown in Tabs.˜13, 13 and 13. As in Amaduzzi et al. (2024; 2025), the reported metrics are a measure of the similarity between the global embeddings of the generated and ground-truth texts, where such embeddings are produced by pre-trained Sentence-BERT (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021) encoders, respectively. Extended results including traditional n-gram-based metrics are shown in App.˜F. Two main conclusions can be drawn from Tabs.˜13, 13, 13, 13, 13 and 13: in the multi-architecture setting, our method is remarkably robust to the choice of test-time NeRF architecture; when instead restricted to MLP-based NeRFs only, our method achieves results comparable to those of the original LLaNA framework. These results further confirm the architecture-agnostic nature of our embeddings and their effectiveness in downstream tasks, even as complex as those involving multimodal NeRF–language understanding.
Table 8: NeRF brief captioning (multi-architecture).
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 67.1 68.9
TRI 66.4 67.8
HASH 66.3 67.9
Table 9: NeRF detailed captioning (multi-architecture).
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 73.3 75.1
TRI 72.2 73.8
HASH 72.4 74.0
Table 10: NeRF single-round Q&A (multi-architecture).
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 80.9 81.4
TRI 80.7 81.3
HASH 80.6 81.3
Table 11: NeRF brief captioning (single-architecture).
Method S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA 68.6 70.5
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 68.3 70.2
Table 12: NeRF detailed captioning (single-architecture).
Method S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA 77.4 79.8
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 74.1 76.0
Table 13: NeRF single-round Q&A (single-architecture).
Method S-BERT $\uparrow$ SimCSE $\uparrow$
LLaNA 81.0 81.6
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 81.0 81.6
@@ 5 Conclusions
In this paper, we presented the first framework that performs downstream tasks on NeRF weights with diverse architectures. Our method, based on a Graph Meta-Network encoder ingesting the NeRF parameter graph, can (i) handle input NeRFs parameterized by MLPs, tri-planes, and, for the first time, multi-resolution hash tables; (ii) process, at inference time, variations of architectures seen during training; (iii) provide state-of-the-art performance in the single-architecture scenario. Furthermore, we investigated the interplay between rendering and contrastive training objectives and showed that they serve complementary purposes, by favoring either class-level separability or invariance to the NeRF architecture. The main limitation of our study lies in its experimental evaluation, which primarily focuses on NeRFs that, although parameterized by 13 total architectures, are trained on a single dataset, i.e. ShapenetRender (Xu et al., 2019), with the notable exception of two generalization experiments on Objaverse (Deitke et al., 2023). Extending our training procedure on Objaverse itself or similar large-scale NeRF datasets would further validate the effectiveness of our approach, and we plan to follow this research path in future work. Eventually, our methodology could be scaled up to become the first foundational model for NeRF weight space processing.
@@ Appendix A Additional Parameter Graph Details
Figure 5: Parameter graph conversion. Top left: parameter graph representation of an MLP, proposed by Lim et al. (2024). Right: parameter graph representation of a tri-plane, proposed by Lim et al. (2024). Dotted edges should be connected to the $C$ channel nodes, but are not fully drawn for better visual clarity. Bottom left: our parameter graph representation of a multi-resolution hash table.
@@
The parameter graph conversion of an MLP, a tri-plane, and a multi-resolution hash table is depicted in Fig.˜5, with additional details compared to Fig.˜2 (left).
@@
MLP. The parameter graph of a linear layer (proposed by Lim et al. (2024)) coincides with its computation graph, except for biases, which are modeled by including an additional node for each (non-input) layer, connected to every other neuron in the layer with an edge. As a result, the parameter graph of an MLP is identical to its computation graph, with the addition of the aforementioned bias nodes. Each node stores three integer values: layer index, neuron index within the layer, and node type (neuron or bias). Each edge stores the corresponding weight or bias and two integers, i.e. layer index and edge type (linear layer weight or linear layer bias).
@@
Tri-plane. The parameter graph of a tri-plane (proposed by Lim et al. (2024)) contains one node for each spatial location and one node for each feature channel, where each spatial node is connected via an edge to every channel node; the channel nodes, together with one node for each input coordinate (more precisely, one node for each entry in their frequency encoding; see App.˜B) become the first layer of the MLP parameter graph that follows. Each node stores three integer values: layer index, neuron index within the layer, and node type (neuron for coordinate nodes or tri-planar for spatial and channel nodes). Each edge stores the corresponding learnable parameter (a.k.a. feature) and five additional values: layer index (integer), edge type (tri-planar, integer), and three $(i,j,k)$ values (evenly spaced floats $\in[-1,1]$) denoting the spatial position of the parameter within the tri-plane (with $i=0$ on the $yz$ plane, $j=0$ in the $xz$ plane, and $k=0$ on the $xy$ plane).
@@
Hash table. As mentioned in Sec.˜3, we are the first to propose a parameter graph construction for multi-resolution hash tables. A single hash table is modeled with a node for each table entry and a node for each feature vector dimension, where each entry node is connected to each feature node via an edge. This subgraph construction is repeated for each hash table, i.e. for each resolution level. Finally, the concatenation of all feature nodes becomes the first layer of the MLP parameter graph that follows. Each node stores three integer values: layer index, neuron index within the layer, and node type (hash table). Each edge stores the corresponding learnable parameter (a.k.a. feature) and four additional integers: layer index, edge type (hash table), table index, and entry index within the table. Consider a multi-resolution hash table with $N$ levels, max table size $T$, $F$ features per entry, and max resolution $R$. Our parameter graph construction requires $N(T+F)$ nodes and $NTF$ edges. Conversely, if we drew an analogy with the tri-planar graph representation and explicitly modeled the underlying voxel grids with one node for each spatial location, it would result in a graph with $\mathcal{O}(N(R^{3}+F))$ nodes and $\mathcal{O}(NR^{3}F)$ edges. Since the purpose of hash grids is to reduce the memory footprint by mapping multiple voxel grid locations to the same table entry via a hashing function, $T$ should always be much smaller than $R^{3}$ (e.g. NeRFs in HASH have $T=2^{12}$ vs $R^{3}=2^{21}$), hence $N(T+F)\ll N(R^{3}+F)$ and $NTF\ll NR^{3}F$. Therefore, the motivation behind our choice of parameter graph representation for hash tables lies in its higher memory efficiency compared to its voxel-grid-based alternative.
@@
Table 14: Dataset architectures. Extended version of Tab.˜1. Architectural hyperparameters of NeRF datasets featured in our experiments, including those used at inference time only (unseen).
@@
Appendix B Additional Experimental Details
NeRF datasets. As outlined in Sec.˜4, the datasets used throughout our experiments mostly consist of NeRFs trained on ShapenetRender (Xu et al., 2019). ShapenetRender is a dataset featuring RGB images and class labels of synthetic objects belonging to 13 classes of the ShapeNetCore dataset (Chang et al., 2015): airplane, bench, cabinet, car, chair, display, lamp, speaker, rifle, sofa, table, phone, and watercraft. For each object, the dataset contains 36 $224\times 224$ images. We follow Zama Ramirez et al. (2024)’s train-val-test split without augmentations, consisting of 31744 NeRFs used for training, 3961 for validation, and 3961 for testing. Tab.˜14 summarizes the NeRF architectures featured in our experiments. In addition to the dataset descriptions provided in Sec.˜4, we note here that:
  • In MLP and TRI, input $(x,y,z)$ coordinates go through a frequency encoding (Mildenhall et al., 2020) before being concatenated to the tri-plane or hash table output that gets fed to the MLP, as this choice was made by the original creators of the two datasets, i.e. Zama Ramirez et al. (2024) and Cardace et al. (2024), respectively. Conversely, the MLP in HASH takes as input the output of the multi-resolution hash table by itself, without (encoded) input coordinates being concatenated to it; we choose to do so in order to be faithful to the original multi-resolution hash table architecture proposed by Müller et al. (2022). The presence or absence of concatenated encoded coordinates in a dataset used for training is then reflected in all the unseen variations of that architecture family (e.g. MLP-2L uses frequency encoding, whereas HASH-2L does not).
  • MLP layers are intertwined with ReLU activation functions in MLP and HASH, whereas TRI uses sine activation functions (Sitzmann et al., 2020). The activation choices for MLP and TRI were made by their original creators (Zama Ramirez et al., 2024; Cardace et al., 2024), whereas we chose ReLU activations for HASH to remain faithful to the original hash table architecture (Müller et al., 2022). The activation choice in a dataset used for training is then reflected in all the unseen variations of that architecture family (e.g. MLP-2L uses ReLU, whereas TRI-2L uses sine).
  • All NeRFs approximate a function from input coordinates to predicted color and density values, i.e. $(x,y,z)\mapsto(r,g,b,\sigma)$. This follows the simplified NeRF formulation proposed by Mildenhall et al. (2020), the alternative being adding a viewing direction $(\theta,\phi)$ as input and letting the NeRF approximate a function $(x,y,z,\theta,\phi)\mapsto(r,g,b,\sigma)$. The formulation with viewing direction is usually modeled by two MLPs: a density MLP taking as input $(x,y,z)$ (or, more precisely, an encoding of $(x,y,z)$) and producing as output $\sigma$ together with an additional vector $\mathbf{h}$ of features, and a color MLP taking as input the concatenation of $\mathbf{h}$ and $(\theta,\phi)$ and predicting $(r,g,b)$ as output. Both Zama Ramirez et al. (2024) and Cardace et al. (2024) adopt the simplified $(x,y,z)\mapsto(r,g,b,\sigma)$ formulation modeled by a single MLP when creating their NeRF datasets (i.e. MLP and TRI, respectively). We choose to follow the same formulation for HASH and the datasets of unseen architectures, both to keep the datasets consistent and to reduce the complexity of the parameter graph going into the encoder.
Furthermore, the experiments in Sec.˜4.3 leverage three additional test sets, dubbed $\texttt{MLP}^{\texttt{OB}}$, $\texttt{TRI}^{\texttt{OB}}$, and $\texttt{HASH}^{\texttt{OB}}$. $\texttt{MLP}^{\texttt{OB}}$ is a dataset of MLP-based NeRFs trained on Objaverse (Deitke et al., 2023), a large-scale dataset of 3D objects. $\texttt{MLP}^{\texttt{OB}}$ was recently released by Amaduzzi et al. (2025), whereas we created $\texttt{TRI}^{\texttt{OB}}$ and $\texttt{HASH}^{\texttt{OB}}$ ourselves by training tri-planar and hash-based NeRFs on the Objaverse test split used in Amaduzzi et al. (2025). More precisely, our test split is the union of the two test splits featured in Amaduzzi et al. (2025)’s benchmark, i.e. the one adopted by PointLLM (Xu et al., 2024) and the one adopted by GPT4Point (Qi et al., 2024). The resulting test set contains 1884 NeRFs. NeRFs in $\texttt{MLP}^{\texttt{OB}}$, $\texttt{TRI}^{\texttt{OB}}$, and $\texttt{HASH}^{\texttt{OB}}$ have the same architecture as those in MLP, TRI, and HASH, respectively.